# Robust Inference for Non-Linear Regression Models with Applications in Enzyme Kinetics

### v1

@@ 6.3 Tyrosinase enzyme data
Both our previous examples had strong outlier effects on estimation and/or testing results. In this final example, we illustrate the performances of the proposed MDPDE based procedures as compared to its competitors for the cases where there is no outlier or only some mild outliers. We consider data from Marasovic et al. [28] which contain two sets of observations on reaction rates and concentrations of the tyrosinase ubiquitous enzyme extracted form two species of mushroom, viz Agaricus bisporus (Ab) and Pleurotus ostreatus (Po). The first set of observations on Ab contain no outlier while there is a mild outlier in the second set on Po. We fitted the MM model (28) to these two datasets using the proposed MDPDE and different competitors. Table 9 contains the parameter estimates, their standard errors and $20\%$ TAPEs; the model fits are shown in Figure 6 along with the scatter plots of the data.
The IWJ method was used for selection of optimum tuning parameter value for these datasets. For Agaricus bisporus (Ab) it took 5 iterations to converge to the optimum value $\widehat{\alpha}=0.35$, while for Pleurotus ostreatus (Po) data the method required only one iteration to select the same optimum value $\widehat{\alpha}=0.53$ for all pilot choices of $\alpha$.
@@
Table 9: Parameter estimates of the MM model fitted to the Tyrosinase enzyme data, along with their asymptotic standard errors (in parenthesis) and $20\%$ TAPE. [For MDPDE, $\widehat{\alpha}=0.35$ and $0.53$ are the optimum choice for Ab and Po data, respectively, obtained by IWJ method]
@@
The scatter plot of Ab data is provided in the Figure 6(a) along with the MDPDE fits (for various $\alpha\geq 0$) and its robust competitors. The figure shows that the proposed MDPDE (for all $\alpha\geq 0$), as well as all robust competitors provide good fit. However, Figure 6(b), corresponding to the Po dataset, shows that the MDPDEs with $\alpha\geq 0.53$ and KPS with $w=1$ provide good robust fits with significantly lower TAPE, than other robust competitors.
We have also tested the validity of the MM model for these data using the Wald-type tests, which resulted a p-value of $<0.001$ for all values of $\alpha\geq 0$ (including the classical Wald test). It shows that the assumed MM model is valid and all tests give the same inference in the presence of no or mild contaminations.

### v2

@@ D.1 Tyrosinase enzyme data
In this example, we illustrate the performance of the proposed MDPDE-based procedure as compared to its competitors for the cases where there is no outlier or only some mild outliers. We consider data from Marasovic et al. [31] which contains observations on reaction rates and concentrations of the tyrosinase ubiquitous enzyme extracted from two species of mushroom, viz Agaricus bisporus (Ab) and Pleurotus ostreatus (Po). The reaction rates of Po are considered in $10^{-4}$. The first set of observations on Ab contains no outlier, while there is a mild outlier in the second set on Po. We fitted the MM model (Eq. (28) of the main paper) to these two datasets using the proposed MDPDE and different competitors. Table 12 contains the parameter estimates, their standard errors, and $20\%$ TAPEs; the model fits are shown in Figure 9 along with the scatter plots of the data.
@@
Table 12: Parameter estimates of the MM model fitted to the Tyrosinase enzyme data, along with their asymptotic standard errors (in parentheses) and $20\%$ TAPE. [For MDPDE, $\widehat{\alpha}=0.35$ and $0.53$ are the optimum choice for Ab and Po data, respectively, obtained by IWJ method]